OpenStudy (anonymous):

Solve for x. x^2+18=6x

OpenStudy (anonymous):

That becomes: \[x ^{2} - 6x + 18 = 0\] Then solve for x using this equation: \[x = \frac{b \pm \sqrt{b^2-4ac}}{2a}\] So, \[x = \frac{-6 \pm \sqrt{(-6)^2-4(1)(18)}}{2(1)}\] But the value enclosed in the square root is negative. This makes it an imaginary number. You will end up with: \[\frac{-6\pm6i}{2}\]